Shawn and Mei are selling pies for a school fundraiser. Customers can buy cherry pies and lemon meringue pies. Shawn sold 10 cherry pies and 13 lemon meringue pies for a total of $282. Mei sold 5 cherry pies and 6 lemon meringue pies for a total of $134. Find the cost of each type of pie.
Answer by ikleyn(52777)

.

From the condition, write the equations as you read


    10C + 13L = 282     (1)

     5C +  6L = 134     (2)


where C, of course, is the number of Cherry pies and L is the number of Lemons.


To solve, multiply second equation by 2, while keeping the first equation as is


    10C + 13L = 282     (1)

    10C + 12L = 268.    (2')


Next, subtract equation (2') from equation (1).  You will get


          13L - 12L = 282 - 268

              L     =    14.


Finally, from equation (2) find C

    5C + 6*14 = 134

    5C +   84 = 134

    5C        = 134 - 84 = 50.

     C                   = 50/5 = 10.


ANSWER.  Cherry pie price is 10 dollars per pie;  Lemon is 14 dollars per pie.

Solved.

------------

The method which I used in my solution is called the ELIMINATION method.